# A model-independent measurement of the expansion and growth rates from BOSS using the FreePower method

### v2

@@ 4 Comparison with previous analyses of BOSS data
A comparison of the inferred cosmological parameters with these works is given in Table 4. In this Table we quote mean values and 68% percentiles using equal-tailed statistics, instead of highest-density values, since those are the ones reported in ISZ20 and PI22. For the low-$z$ survey we find $f=0.66\pm 0.12$ and $E=1.07\pm 0.06$, in reasonable agreement with $f=0.74\pm 0.08$, $E=1.14\pm 0.05$ from ISZ20. FreePower measures directly $f(z)$, while conventional analysis measure instead $f(z)\sigma_{8}(z)$. We therefore convert the latter using the values of $\sigma_{8}$ and $\Omega_{m0}$ obtained in ISZ20 and propagate the uncertainty in the integral for the growth factor.
@@
par. this work ISZ20 PI22 this work ISZ20 PI22
$z_{\rm eff}=0.38$ $z_{\rm eff}=0.61$
$f$ $0.66\pm 0.12$ $0.735\pm 0.084$ $0.64\pm 0.14$ $0.750\pm 0.085$
$E$ $1.07\pm 0.06$ $1.136\pm 0.046$ $1.29\pm 0.11$ $1.384\pm 0.073$
$b_{1}$ $2.14\pm 0.18$ $1.67^{+0.16}_{-0.20}$ $2.21\pm 0.14$ $2.16^{+0.21}_{-0.22}$ $2.28^{+0.22}_{-0.30}$ $2.33\pm 0.15$
$b_{2}$ $-1.6^{+1.6}_{-1.7}$ $-2.3^{+0.5}_{-0.9}$ $-0.51^{+0.79}_{-0.97}$ $-0.55^{+2.84}_{-2.54}$ $-2.7^{+0.5}_{-2.1}$ $-1.09^{+0.84}_{-1.0}$
$b_{\mathcal{G}_{2}}$ $0.35^{+0.35}_{-0.32}$ $0.11^{+0.28}_{-0.40}$ $-0.38\pm 0.37$ $-0.21\pm 0.78$ $0.59^{+0.26}_{-0.08}$ $-0.19\pm 0.43$
Our errors are in many cases comparable, but in some cases such as for $f$ and $E$ at $z_{\rm eff}=0.61$ they are between 30 and 50% larger than previous analyses based on $\Lambda$CDM. A smaller precision is of course entirely expected since our approach is more flexible and introduces more free parameters. We nevertheless find it noteworthy that the loss of precision is in general not very large, especially compared with the gains in robustness to modelling.
@@ 6 Conclusions and outlook
  • We were able to constrain both non-cosmological parameters (bias, shot-noise and counterterm parameters), and cosmological parameters in a model-independent way using LSS surveys. In particular, we measured $f(z)$ and $E(z)$ at both BOSS redshift bins (Table 2).
  • Our constraints are consistent with the ones from previous model-dependent analyses [14, 17]. In the case of the growth rate and the expansion rate, the uncertainties increase in the model-independent analysis, as expected due to it having fewer assumptions.
  • For $z_{\rm eff}=0.38$, we find a tension at over $2\sigma$ with $\Lambda$CDM results for $E$, but not for the growth rate $f$. For $z_{\rm eff}=0.61$ we find good agreement with $\Lambda$CDM for both $E$ and $f$. For the bias parameters we find good agreement with a previous 1-loop analysis in both bins.
@@
There are of course many ways to extend our results in the future. In terms of data, the hexadecapole can be included in the analysis next to monopole and quadrupole. One can also include the bispectrum [56], which is of the same order as the one-loop power spectrum in the non-linear perturbation theory [17]. The BOSS SGC data can also be included, which would increase the effective volume of the sample by around 37% [14]. In terms of methodology, the growth rate $f(z)$ could be generalized to be scale-dependent, and/or one could try using a larger number of $k$-bins. One could also use uninformative priors for the dimensionless distance $L_{A}$ to let it be constrained solely by the LSS data. Another possibility is to perform a combined analysis of both redshift bins, since in FreePower they share the same linear $P(k)$ parameters. In terms of theory, our analysis can be expanded by including deviations from the EdS kernels [57, 58, 59]. It can also be extended to include the peculiar-velocity power spectrum [27, 60, 61]. We note that a similar model-independent analysis can be carried out to investigate possible primordial non-Gaussianity (PNG) [16]. We plan to pursue these lines of investigation in future work.
@@ Acknowledgements
We thank Sandro Vitenti for help with the MCMC convergence analysis, Massimo Pietroni and Marco Marinucci for several discussions and Eoin Colgáin for comments on our manuscript. MQ is supported by the Brazilian research agencies Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) project E-26/201.237/2022, CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) and CAPES. LA acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster) and under project 456622116. We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E-26/210.130/2023. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. We acknowledge support from the CAPES-DAAD bilateral project “Data Analysis and Model Testing in the Era of Precision Cosmology”.
@@ Appendix A Correlations and convergence
The correlation matrices among all 17 parameters are depicted in Figure 6, for both redshift bins. As discussed in the text, most terms are small. The bias parameters have stronger correlations among themselves for the lower redshift bin. The different $P(k)$ bins have strong positive correlations with their immediate neighbors, with the exception of $P_{1}$, for which tight priors were used. The data-driven cosmological parameters $f$ and $E$ are positively correlated among themselves and anti-correlated with $b_{1}$ and $b_{\mathcal{G}_{2}}$.
We also display the integrated autocorrelation times ($\tau$) in Figure 7. We compute $\tau$ for each walker individually and these estimates are averaged afterwards [37]. The MCMC shows clear signs of convergence as the estimates for the $\tau$ start crossing the $N/20$ line, where $N$ is the number of MCMC steps. Note that since $\tau$ is of the order $10^{4}$ for all parameters, we thin all our chains by a factor of 100 in our analysis, which cuts some of the noise and allows more robust KDE estimation and thus more robust confidence intervals with getdist.
Figure 7: Integrated autocorrelation times as a function of the number of steps in each of the 256 walkers. Top: $z_{\rm eff}=0.38$; bottom: $z_{\rm eff}=0.61$.
@@ Appendix C One-loop galaxy power spectrum
The one-loop corrections for the galaxy density field power spectrum take the form [14]
@@
Figure 9: Same as Figure 8 for $z_{\rm eff}=0.61$.
Figure 11: Same as Figure 10 for $z_{\rm eff}=0.61$.
The kernels have been derived for an Einstein-deSitter cosmology. They include three free $z$-dependent functions, denoted $b_{1}$, $b_{2}$, $b_{\mathcal{G}_{2}}$. The bias parameter $b_{\Gamma_{3}}$ is set to zero. The restriction to Einstein-deSitter is of course a break of the model-independence, although it is well-known that this approximation works very well also for $\Lambda$CDM and other cosmologies that do not depart too much from it. In our FreePower method the kernels have been implemented in a much more general way, under general conditions like equivalence principle and Galilean invariance, and can be considered essentially model-independent.
The counterterms can be parametrized as follows [62]:

### v3

@@ 4 Comparison with previous analyses of BOSS data
A comparison of the inferred cosmological parameters with these works is given in Table 4. In this Table we quote mean values and 68% percentiles using equal-tailed statistics, instead of highest-density values, since those are the ones reported in ISZ20 and PI22. For the low-$z$ survey we find $f=0.7\pm 0.2$ and $E=1.08\pm 0.10$, in reasonable agreement with $f=0.74\pm 0.08$, $E=1.14\pm 0.05$ from ISZ20. FreePower measures directly $f(z)$, while conventional analysis measure instead $f(z)\sigma_{8}(z)$. We therefore convert the latter using the values of $\sigma_{8}$ and $\Omega_{m0}$ obtained in ISZ20 and propagate the uncertainty in the integral for the growth factor.
par. this work ISZ20 PI22 this work ISZ20 PI22
$z_{\rm eff}=0.38$ $z_{\rm eff}=0.61$
$f$ $0.70^{+0.20}_{-0.19}$ $0.735\pm 0.084$ $0.87^{+0.23}_{-0.22}$ $0.750\pm 0.085$
$E$ $1.08^{+0.10}_{-0.11}$ $1.138\pm 0.049$ $1.15\pm 0.11$ $1.386\pm 0.076$
$L_{A}$ $0.25^{+0.03}_{-0.04}$ $0.242\pm 0.006$ $0.38\pm 0.04$ $0.309\pm 0.012$
$b_{1}$ $2.16\pm 0.33$ $1.67^{+0.16}_{-0.20}$ $2.21\pm 0.14$ $2.57\pm 0.36$ $2.28^{+0.22}_{-0.30}$ $2.33\pm 0.15$
$b_{2}$ $-1.48^{+1.70}_{-1.86}$ $-2.3^{+0.5}_{-0.9}$ $-0.51^{+0.79}_{-0.97}$ $-0.34^{+3.45}_{-3.14}$ $-2.7^{+0.5}_{-2.1}$ $-1.09^{+0.84}_{-1.00}$
$b_{\mathcal{G}_{2}}$ $0.34^{+0.44}_{-0.39}$ $0.11^{+0.28}_{-0.40}$ $-0.38\pm 0.37$ $0.12\pm 0.85$ $0.59^{+0.26}_{-0.08}$ $-0.19\pm 0.43$
@@
As expected, our errors are always significantly larger than in model-dependent analyses. For instance, the errors on $f$, $E$ and $L_{A}$ are roughly twice as large.
@@ 6 Conclusions and outlook
  • We were able to constrain both non-cosmological parameters (bias, shot-noise and counterterm parameters), and cosmological parameters in a model-independent way using LSS surveys. In particular, we measured $f(z),E(z)$ and $L_{A}(z)$ at both BOSS redshift bins (Table 2).
  • Our constraints are consistent with the ones from previous model-dependent analyses [14, 17]. The uncertainties increase in the model-independent analysis, as expected due to it having fewer assumptions.
[…]
@@
There are of course many ways to extend our results in the future. In terms of data, the hexadecapole can be included in the analysis next to monopole and quadrupole. One can also include the bispectrum [58], which is of the same order as the one-loop power spectrum in the non-linear perturbation theory [17]. The BOSS SGC data can also be included, which would increase the effective volume of the sample by around 37% [14]. In terms of methodology, the growth rate $f(z)$ could be generalized to be scale-dependent, and/or one could try using a larger number of $k$-bins. One could also use more informative priors for the dimensionless distance $L_{A}$ taken from external datasets like SN Ia. Another possibility is to perform a combined analysis of both redshift bins, since in FreePower they share the same linear $P(k)$ parameters. In terms of theory, our analysis can be expanded by including deviations from the EdS kernels [59, 60, 61]. It can also be extended to include the peculiar-velocity power spectrum [27, 62, 63]. We note that a similar model-independent analysis can be carried out to investigate possible primordial non-Gaussianity (PNG) [16]. We plan to pursue these lines of investigation in future work.
@@ Acknowledgements
We thank Sandro Vitenti for help with the MCMC convergence analysis, Louis Legrand and Vivian Miranda for tips regarding MCMC samplers, Massimo Pietroni and Marco Marinucci for several discussions and Eoin Colgáin for comments on our manuscript. MQ is supported by the Brazilian research agencies Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) project E-26/201.237/2022, CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) and CAPES. LA acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster) and under project 456622116. We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E-26/210.130/2023. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. We acknowledge support from the CAPES-DAAD bilateral project “Data Analysis and Model Testing in the Era of Precision Cosmology”.
@@ Appendix A Correlations and convergence
The correlation matrices among all 17 parameters are depicted in Figure 6, for both redshift bins. As discussed in the text, most terms are small. The bias parameters have stronger correlations among themselves for the lower redshift bin. The different $P(k)$ bins have strong positive correlations with their immediate neighbors, with the exception of $P_{1}$ and $P_{2}$, for which tight priors were used. The data-driven cosmological parameters $f$, $E$ and $L_{A}$ are correlated among themselves and $b_{1}$. $E$ is also correlated with $b_{\mathcal{G}_{2}}$.
We also display the integrated autocorrelation times ($\tau$) in Figure 7. We compute $\tau$ for each walker individually and these estimates are averaged afterwards [47]. The MCMC shows clear signs of convergence as the estimates for the $\tau$ crosses the $N/50$ line, where $N$ is the number of MCMC steps. Note that $\tau\sim 10$ for all parameters, but we ran many more than 500 steps to get smoother 2D contours. We do not perform any thinning of the chains.
Figure 7: Integrated autocorrelation times as a function of the number of steps in each of the 60 walkers. Top: $z_{\rm eff}=0.38$; bottom: $z_{\rm eff}=0.61$.
Regarding the Gelman Rubin test, for all parameters we get $(\hat{R})^{(1/2)}-1<0.0001$ for both low and high-$z$ runs. We note that $(\hat{R})^{(1/2)}$ is not the most reliable test specially using ensemble samplers like ptemcee, since in this approach the walkers are not independent. Nevertheless, since the results are really very close to unity, they also point to a very good convergence of all chains.
@@ Appendix C One-loop galaxy power spectrum
Figure 9: Same as Figure 8 for $z_{\rm eff}=0.61$.
Figure 11: Same as Figure 10 for $z_{\rm eff}=0.61$.
The one-loop corrections for the galaxy density field power spectrum have studied in many works (see e.g. the review [64]) and, when applied to tracers, take the form [14]
@@
The kernels we adopted in this work have been derived for an Einstein-deSitter cosmology. They include three free $z$-dependent functions, denoted $b_{1}$, $b_{2}$, $b_{\mathcal{G}_{2}}$. The bias parameter $b_{\Gamma_{3}}$ is set to zero. The restriction to Einstein-deSitter is of course a break of the model-independence, although it is well-known that this approximation works very well also for $\Lambda$CDM and other cosmologies that do not depart too much from it. In our FreePower method the kernels have been implemented in a much more general way, under general conditions like equivalence principle and Galilean invariance, and can be considered essentially model-independent.
The counterterms are added to take into account the rotational part of the flow and improve convergence. They can be parametrized as follows [65]: